# SoloSpeech: Enhancing Intelligibility and Quality in Target Speech Extraction through a Cascaded Generative Pipeline

### v2

@@ Abstract
Target Speech Extraction (TSE) aims to isolate a target speaker’s voice from a mixture of multiple speakers by leveraging speaker-specific cues, typically provided as auxiliary audio (a.k.a. cue audio). Although recent advancements in TSE have primarily employed discriminative models that offer high perceptual quality, these models often introduce unwanted artifacts, reduce naturalness, and are sensitive to discrepancies between training and testing environments. On the other hand, generative models for TSE lag in perceptual quality and intelligibility. To address these challenges, we present SoloSpeech, a novel cascaded generative pipeline that integrates compression, extraction, reconstruction, and correction processes. SoloSpeech features a speaker-embedding-free target extractor that utilizes conditional information from the cue audio’s latent space, aligning it with the mixture audio’s latent space to prevent mismatches. Evaluated on the widely-used Libri2Mix dataset, SoloSpeech achieves the new state-of-the-art intelligibility and quality in target speech extraction and speech separation tasks while demonstrating exceptional generalization on out-of-domain data and real-world scenarios.
1 Introduction
@@
Humans possess the remarkable ability to focus on a specific speaker amidst a cacophony of sounds, including background noise, reverberation, and interfering voices. This capability is famously referred to as the cocktail-party effect bronkhorst2015cocktail . For many years, researchers have endeavored to emulate this sophisticated auditory skill using algorithms. Target Speech Extraction (TSE) emerges as a promising field, aiming to isolate a desired speaker’s voice from a complex mixture of multiple speakers, alongside any accompanying noise or reverberation DBLP:journals/spm/ZmolikovaDOKCY23 . TSE leverages cues, which serve as identifiers for the target speaker, to accurately extract their speech from a mixture. Various types of cues have been investigated to address the TSE problem. These include spatial cues DBLP:conf/interspeech/GuCZZXYSZ019 ; DBLP:conf/icassp/SubramanianWYZX20 , visual cues DBLP:journals/tog/EphratMLDWHFR18 , concept cues DBLP:conf/mm/OhishiDOATNKHK22 , and audio cues DBLP:conf/interspeech/WangMWSWHSWJL19 ; DBLP:journals/jstsp/ZmolikovaDKONBC19 . In this paper, we specifically focus on monaural TSE methods that utilize the most prevalent audio-based cues.
@@
A significant challenge in TSE lies in distinguishing between target and non-target speakers, as their speech signals often share similar characteristics, complicating the extraction process DBLP:conf/interspeech/ZhaoYGZZ22 . To overcome this, many researchers have concentrated on discriminative model-based TSE, which directly maps the mixture signal to the target signal by optimizing signal-level metrics han2022dpccn ; DBLP:conf/interspeech/WangYWYZ22 ; DBLP:journals/corr/abs-2409-15799 . Both time-domain DBLP:journals/taslp/XuRCL20 ; DBLP:conf/interspeech/GeXWCD020 ; DBLP:conf/icassp/LiuDWZ23 and time-frequency (T-F) domain methodologies DBLP:journals/inffus/HaoLZ24 ; DBLP:journals/corr/abs-2409-02615 have been investigated. While these discriminative approaches are highly effective, they can sometimes introduce artifacts and are often sensitive to discrepancies between training and testing conditions kamo23_interspeech ; wang2024noise . In contrast, generative approaches model the joint distribution of the mixture speech, the target speech, and the cue speech. Various deep generative models have been explored for tasks such as speech enhancement, dereverberation, and separation DBLP:journals/taslp/RichterWLLG23 ; scheibler2023diffusion ; DBLP:conf/iclr/Liu0VSTH24 . These models have demonstrated the ability to exhibit greater robustness in unseen scenarios than discriminative models DBLP:journals/taslp/RichterWLLG23 ; wang2024noise . They are capable of achieving high clarity in overlapping regions where multiple speakers converse simultaneously DBLP:journals/corr/abs-2409-08425 . However, current generative approaches still result in separated speech with reduced audio quality and intelligibility, leading mainstream TSE solutions to continue relying on discriminative methods DBLP:journals/corr/abs-2301-10752 .

### v3

@@ Abstract
Target Speech Extraction (TSE) aims to isolate a target speaker’s voice from a mixture of multiple speakers by leveraging speaker-specific cues, typically provided as auxiliary audio (a.k.a. cue audio). Although recent advancements in TSE have primarily employed discriminative models that offer high perceptual quality, these models often introduce unwanted artifacts, reduce naturalness, and are sensitive to discrepancies between training and testing environments. On the other hand, generative models for TSE lag in perceptual quality and intelligibility. To address these challenges, we present SoloSpeech, a novel cascaded generative pipeline that integrates compression, extraction, reconstruction, and correction processes. SoloSpeech features a speaker-embedding-free target extractor that utilizes conditional information from the cue audio’s latent space, aligning it with the mixture audio’s latent space to prevent mismatches. Evaluated on the widely-used Libri2Mix dataset, SoloSpeech achieves the new state-of-the-art intelligibility and quality in TSE while demonstrating exceptional generalization on out-of-domain data and real-world scenarios.
Index Terms:
Target speech extraction, cocktail-party effect, generative model, latent diffusion model, out-of-domain data.
I Introduction
@@
Humans possess the remarkable ability to focus on a specific speaker amidst a cacophony of sounds, including background noise, reverberation, and interfering voices. This capability is famously referred to as the cocktail-party effect [1]. For many years, researchers have endeavored to emulate this sophisticated auditory skill using algorithms. Target Speech Extraction (TSE) emerges as a promising field, aiming to isolate a desired speaker’s voice from a complex mixture of multiple speakers, alongside any accompanying noise or reverberation [2]. TSE leverages cues, which serve as identifiers for the target speaker, to accurately extract their speech from a mixture. Various types of cues have been investigated to address the TSE problem. These include spatial cues [3, 4], visual cues [5], concept cues [6], and audio cues [7, 8, 9]. In this wor, we specifically focus on monaural TSE methods that utilize the most prevalent audio-based cues.
@@
A significant challenge in TSE lies in distinguishing between target and non-target speakers, as their speech signals often share similar characteristics, complicating the extraction process [10]. To overcome this, many researchers have concentrated on discriminative model-based TSE, which directly maps the mixture signal to the target signal by optimizing signal-level metrics [11, 12, 13]. Both time-domain [14, 15, 16] and time-frequency (T-F) domain methodologies [17, 18] have been investigated. While these discriminative approaches are highly effective, they can sometimes introduce artifacts and are often sensitive to discrepancies between training and testing conditions [19, 20]. In contrast, generative approaches model the joint distribution of the mixture speech, the target speech, and the cue speech. Various deep generative models have been explored for tasks such as speech enhancement, dereverberation, and separation [21, 22, 23, 24]. These models have demonstrated the ability to exhibit greater robustness in unseen scenarios than discriminative models [21, 20]. They are capable of achieving high clarity in overlapping regions where multiple speakers converse simultaneously [25]. However, current generative approaches still result in separated speech with reduced audio quality and intelligibility, leading mainstream TSE solutions to continue relying on discriminative methods [26].